An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 786 hours and a standard deviation of 80 hours. Find the probability that a random samples of 17 bulbs will have an average life between 799 and 802 hours.


1
Expert's answer
2022-02-07T16:17:09-0500

We use Z calculations

z(799) = (799 - 786)/80 = 0.1625

z(802) = ( 802 - 786)/80 = 0.2

We then define

P (0.1625 < z < 0.2), using the normal table we have

( 0.57926 - 0.56356) = 0.01561 which is the required probability.